# Optimal Computation in Anonymous Dynamic Networks

### v7

@@ 3.3 Related Concepts
In the history tree of a 1-union-connected dynamic network, every node at level $L_{t}$ is in the view of every node at level $L_{t^{\prime}}$, for all $t^{\prime}\geq t+n-1$.
@@
Let $v\in L_{t}$, and let $P$ be the set of processes not represented by $v$. If $P$ is empty, then all nodes in $L_{t^{\prime}}$ are descendants of $v$, and have $v$ in their view. Otherwise, $11\leq|P|\leq n-$ 1, and Lemma 3.3 implies that $v$ is in the view of all nodes in $L_{t^{\prime}}$. ∎
@@ 4 Leaderless and Stabilizing Algorithms
Leveraging the theory developed in Section 3.2, we will assume, without any loss of generality, that all processes maintain their current view of the history tree as their internal state and broadcast this view across all available links at every round.
@@ 4.1 Basic Technique
Our basic technique makes use of the procedure in Listing 1, which will be a subroutine of all algorithms presented in this section. The purpose of this procedure is to construct a homogeneous system [^20] A linear system is homogeneous if all its constant terms are zero. of $k-1$ independent linear equations involving the anonymities of all the $k$ nodes in a level of a process’ view. We will first give some definitions.
@@
Exposed nodes and strands. In (a view of) a history tree, if a node $v\in L_{t}$ has exactly one child (i.e., there is exactly one node $v^{\prime}\in L_{t+1}$ such that $\{v,v^{\prime}\}$ is a black edge), we say that $v$ is non-branching. If $v$ has multiple children, it is branching (thus, a leaf in a view is neither branching nor non-branching). We emphasize that, if $v$ is a node of a view $\mathcal{V}$ of a history tree $\mathcal{H}$, then $v$ may be branching in $\mathcal{H}$ but non-branching in $\mathcal{V}$. This happens if $v$ has multiple children in $\mathcal{H}$, but only one of them appears in $\mathcal{V}$.
We say that two non-branching nodes $v_{1},v_{2}\in L_{t}$, whose respective children are $v^{\prime}_{1},v^{\prime}_{2}\in L_{t+1}$, are exposed with multiplicity $(m_{1},m_{2})$ if the red edges $\{v^{\prime}_{1},v_{2}\}$ and $\{v^{\prime}_{2},v_{1}\}$ are present with multiplicities $m_{1}\geq 1$ and $m_{2}\geq 1$, respectively (see Figure 3). Again, the same two nodes may be exposed in a view of a history tree, but not in the history tree itself.
A strand is a path $(w_{1},w_{2},\dots,w_{k})$ in (a view of) a history tree consisting of non-branching nodes such that, for all $1\leq i<k$, the node $w_{i}$ is the parent of $w_{i+1}$. We say that two strands $P_{1}$ and $P_{2}$ are exposed if there are two exposed nodes $v_{1}\in P_{1}$ and $v_{2}\in P_{2}$.
@@
Let $v_{1},v_{2}\in L_{t}$, and let $P_{1}$ and $P_{2}$ be the sets of processes represented by $v_{1}$ and $v_{2}$, respectively. Since $v_{1}$ is non-branching in $\mathcal{H}$, we have $a(c(v_{1}))=a(v_{1})$, and therefore $c(v_{1})$ represents $P_{1}$, as well. Hence, the number of links between $P_{1}$ and $P_{2}$ in $G_{t+1}$ (counted with their multiplicities) is $a(c(v_{1}))\cdot m_{1}=a(v_{1})\cdot m_{1}$. By a symmetric argument, this number is equal to $a(v_{2})\cdot m_{2}$. ∎
Observe that Lemma 4.1 may not hold in a view $\mathcal{V}$ of $\mathcal{H}$, because $v_{1}$ (or $v_{2}$) may be non-branching in $\mathcal{V}$ but have multiple children in $\mathcal{H}$. Thus, it is not necessarily true that $v_{1}$ (or $v_{2}$) and its unique child in $\mathcal{V}$ have the same anonymity.
Main subroutine. Intuitively, the procedure in Listing 1 searches for a long-enough sequence of levels in the given view $\mathcal{V}$, say from $L_{s}$ to $L_{t}$, where all nodes are non-branching. That is, the nodes in $L_{s}\cup L_{s+1}\cup\dots\cup L_{t}$ can be partitioned into $k=|L_{s}|=|L_{t}|$ strands. Then the procedure searches for pairs of exposed strands, each of which yields a linear equation involving the anonymities of some nodes of $L_{t}$, until it obtains $k-1$ linearly independent equations. [^21] The reason why we have to consider strands spanning several levels of the history tree (as opposed to looking at a single level) is that the dynamic disconnectivity $\tau$ is not known, and thus Proposition 2.4 cannot be applied directly. Note that the search may fail (in which case Listing 1 returns $t=-1$) or it may produce incorrect equations. The following lemma specifies sufficient conditions for Listing 1 to return a correct and non-trivial system of equations for some $t\geq 0$.
Listing 1: Constructing a system of equations in the anonymities of some nodes in a view.
1# Input: a view $\mathcal{V}$ with levels $L_{-1}$, $L_{0}$, $L_{1}$, $\dots$, $L_{h}$
@@ 20 Return $(t,S)$
Let $\mathcal{V}$ be the view of a process in a $\tau$-union-connected network of size $n$ taken at round $t^{\prime}$, and let Listing 1 return $(t,S)$ on input $\mathcal{V}$. Assume that one of the following conditions holds:
@@
    $t^{\prime}\geq\tau(2n-c-1)+1$, where $c$ is the number of distinct inputs held by processes at round 00.
@@
If $\tau=1$, it takes at most $n-1$ rounds for information to travel from a process to any other process, due to Corollary 3.4. In general, if $\tau\geq 1$, it takes at most $\tau(n-1)$ rounds, by Proposition 2.4. Therefore, since $\mathcal{V}$ is a view taken at round $t^{\prime}$, all levels of $\mathcal{V}$ up to $L_{t^{\prime}-\tau(n-1)}$ are complete (recall that a level of a view is complete if it has no missing nodes).
Assume Condition 2 first. Since $t^{\prime}\geq\tau(2n-c-1)+1$, all levels of $\mathcal{V}$ up to $L_{\tau(n-c)+1}$ are complete. Thus, level $L_{0}$ is complete, and therefore it has exactly $c$ nodes, because two processes are distinguishable at round 00 if and only if they have distinct inputs. Since the sum of the anonymities of these $c$ nodes is $n$, there must be fewer than $n-c$ branching nodes in $\mathcal{V}$ (excluding the root). Hence, the levels up to $L_{\tau(n-c)}$ contain an interval of at least $\tau$ consecutive levels, say from $L_{r}$ to $L_{r+\tau-1}$, where all nodes are non-branching and can be partitioned into $|L_{r}|=|L_{r+\tau-1}|$ strands $P_{1}$, $P_{2}$, …, $P_{|L_{r}|}$.
Note that a link between two processes at any round $r^{\prime}$ in the interval $[r+1,r+\tau]$ determines a pair of exposed nodes in $L_{r^{\prime}-1}$. Thus, by definition of $\tau$-union-connected network, the graph of exposed strands between $L_{r}$ and $L_{r+\tau-1}$ (constructed as $G$ in Line 12) is connected. It follows that the execution of Listing 1 terminates at Line 20 (as opposed to Line 6) whenever $t\geq r+\tau-1$. Thus, the procedure returns a pair $(t,S)$ with $0\leq t\leq r+\tau-1\leq\tau(n-c)$. In particular, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete.
@@
    $t^{\prime}\geq\tau(n-c)+d+1$, where $c$ is the number of distinct inputs held by processes at round 00,
@@ 4.2 Stabilizing Algorithm for Leaderless Networks
As a first application of Lemma 4.2, we will give stabilizing algorithm that efficiently computes the Input Frequency function $F_{IF}$ in all leaderless networks with a finite dynamic disconnectivity $\tau$. As a consequence, all frequency-based functions are efficiently computable as well, due to Proposition 2.2. Moreover, Proposition 6.3 states that no other functions are computable in leaderless networks, and Proposition 6.9 shows that our algorithm is asymptotically optimal.
@@
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and stabilizes in at most $\tau(2n-2)$ rounds, assuming no knowledge of $\tau$ or $n$.
@@
Our local algorithm is as follows. Each process $p$ runs Listing 1 on its own view $\mathcal{V}$, obtaining a pair $(t,S)$. If $t=-1$ or $S$ is not a homogeneous system of $k-1$ independent linear equations in $k$ variables, then $p$ outputs $\{(\lambda(p),1)\}$, where $\lambda(p)$ is its own input. Otherwise, since the rank of the coefficient matrix of $S$ is $k-1$, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (recall that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, $p$ outputs
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 4.2. Specifically, if not all processes have the same input (hence $n\geq c\geq 2$), then at any round $t^{\prime}\geq\tau(2n-2)\geq\tau(2n-3)+1\geq\tau(2n-c-1)+1$, Condition 2 of Lemma 4.2 is met, and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $\tau(2n-2)$ rounds.
In the special case where all processes have the same input, then the default output $\{(\lambda(p),1)\}$ is correct since round 00. Also, whenever $S$ is a homogeneous system of $k-1$ independent linear equations, the algorithm still returns the correct output $\{(\lambda(p),1)\}$, because all nodes in level $L_{0}$ have the same label $\lambda(p)$. Hence, in this particular case the output is always correct. ∎
@@ 4.3 Terminating Algorithm for Leaderless Networks
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and terminates in at most $\tau(n+N-2)+1$ rounds, assuming that $\tau$ and an upper bound $N\geq n$ are known to all processes.
@@
Assume that $\tau$ and $N$ are known. Our terminating local algorithm is as follows. Run Listing 1 on the process’ view $\mathcal{V}$, obtaining a pair $(t,S)$, and then do the same computations as in the algorithm for Theorem 4.4. If $t\geq 0$ and the current round $t^{\prime}$ satisfies $t^{\prime}\geq t+\tau(N-1)+1$, then the output is correct, and the process terminates.
@@
We can also trade the knowledge of $\tau$ and $N$ for the knowledge of the dynamic diameter $d$ of the network.
@@
There is an algorithm that computes $F_{IF}$ in all $\tau$-union-connected anonymous networks with no leader and terminates in at most $\tau(n-1)+d+1\leq dn+1$ rounds, assuming that the dynamic diameter (or an upper bound thereof) $d$ is known to all processes, but assuming no knowledge of $\tau$ or $n$.

### v8

@@ 3.3 Related Concepts
In the history tree of a 1-union-connected dynamic network, every node at level $L_{t}$ is in the vista of every node at level $L_{t^{\prime}}$, for all $t^{\prime}\geq t+n-1$.
@@
Let $v\in L_{t}$, and let $P$ be the set of agents not represented by $v$. If $P$ is empty, then all nodes in $L_{t^{\prime}}$ are descendants of $v$, and have $v$ in their vista. Otherwise, $11\leq|P|\leq n-$ 1, and Lemma 3.3 implies that $v$ is in the vista of all nodes in $L_{t^{\prime}}$. ∎
@@ 4 Leaderless and Stabilizing Algorithms
Leveraging the theory developed in Section 3.2, we will assume, without any loss of generality, that all agents maintain their current vista of the history tree as their internal state and broadcast this vista across all available links at every round.
@@ 4.1 Basic Technique
Our basic technique makes use of the procedure in Listing 1, which will be a subroutine of all algorithms presented in this section. The purpose of this procedure is to construct a homogeneous system of $k-1$ independent linear equations involving the anonymities of all the $k$ nodes in a level of an agent’s vista (recall that a linear system is homogeneous if all its constant terms are zero). We will first give some definitions.
@@
Exposed nodes and strands. In (a vista of) a history tree, if a node $v\in L_{t}$ has exactly one child (i.e., there is exactly one node $v^{\prime}\in L_{t+1}$ such that $\{v,v^{\prime}\}$ is a black edge), we say that $v$ is non-branching. If $v$ has multiple children, it is branching (thus, a leaf in a vista is neither branching nor non-branching). We emphasize that, if $v$ is a node of a vista $\mathcal{V}$ of a history tree $\mathcal{H}$, then $v$ may be branching in $\mathcal{H}$ but non-branching in $\mathcal{V}$. This happens if $v$ has multiple children in $\mathcal{H}$, but only one of them appears in $\mathcal{V}$.
We say that two non-branching nodes $v_{1},v_{2}\in L_{t}$, whose respective children are $v^{\prime}_{1},v^{\prime}_{2}\in L_{t+1}$, are exposed with multiplicity $(m_{1},m_{2})$ if the red edges $\{v^{\prime}_{1},v_{2}\}$ and $\{v^{\prime}_{2},v_{1}\}$ are present with multiplicities $m_{1}\geq 1$ and $m_{2}\geq 1$, respectively (see Figure 3). Again, the same two nodes may be exposed in a vista of a history tree, but not in the history tree itself.
A strand is a path $(w_{1},w_{2},\dots,w_{k})$ in (a vista of) a history tree consisting of non-branching nodes such that, for all $1\leq i<k$, the node $w_{i}$ is the parent of $w_{i+1}$. We say that two strands $P_{1}$ and $P_{2}$ are exposed if there are two exposed nodes $v_{1}\in P_{1}$ and $v_{2}\in P_{2}$.
@@
Let $v_{1},v_{2}\in L_{t}$, and let $P_{1}$ and $P_{2}$ be the sets of agents represented by $v_{1}$ and $v_{2}$, respectively. Since $v_{1}$ is non-branching in $\mathcal{H}$, we have $a(c(v_{1}))=a(v_{1})$, and therefore $c(v_{1})$ represents $P_{1}$, as well. Hence, the number of links between $P_{1}$ and $P_{2}$ in $G_{t+1}$ (counted with their multiplicities) is $a(c(v_{1}))\cdot m_{1}=a(v_{1})\cdot m_{1}$. By a symmetric argument, this number is equal to $a(v_{2})\cdot m_{2}$. ∎
Observe that Lemma 4.1 may not hold in a vista $\mathcal{V}$ of $\mathcal{H}$, because $v_{1}$ (or $v_{2}$) may be non-branching in $\mathcal{V}$ but have multiple children in $\mathcal{H}$. Thus, it is not necessarily true that $v_{1}$ (or $v_{2}$) and its unique child in $\mathcal{V}$ have the same anonymity.
Main subroutine. Intuitively, the procedure in Listing 1 searches for a long-enough sequence of levels in the given vista $\mathcal{V}$, say from $L_{s}$ to $L_{t}$, where all nodes are non-branching. That is, the nodes in $L_{s}\cup L_{s+1}\cup\dots\cup L_{t}$ can be partitioned into $k=|L_{s}|=|L_{t}|$ strands. Then the procedure searches for pairs of exposed strands, each of which yields a linear equation involving the anonymities of some nodes of $L_{t}$, until it obtains $k-1$ linearly independent equations. The reason why we have to consider strands spanning several levels of the history tree (as opposed to looking at a single level) is that the dynamic disconnectivity $\tau$ is not known, and thus Proposition 2.5 cannot be applied directly. Note that the search may fail (in which case Listing 1 returns $t=-1$) or it may produce incorrect equations. The following lemma specifies sufficient conditions for Listing 1 to return a correct and non-trivial system of equations for some $t\geq 0$.
Listing 1: Constructing a system of equations in the anonymities of some nodes in a vista.
1# Input: a vista $\mathcal{V}$ with levels $L_{-1}$, $L_{0}$, $L_{1}$, $\dots$, $L_{h}$
@@ 20 Return $(t,S)$
Let $\mathcal{V}$ be the vista of an agent in a $\tau$-union-connected network of size $n$ taken at round $t^{\prime}$, and let Listing 1 return $(t,S)$ on input $\mathcal{V}$. Assume that one of the following conditions holds:
@@
    $t^{\prime}\geq\tau(2n-c-1)+1$, where $c$ is the number of distinct inputs held by agents at round 00.
@@
If $\tau=1$, it takes at most $n-1$ rounds for information to travel from an agent to any other agent, due to Corollary 3.4. In general, if $\tau\geq 1$, it takes at most $\tau(n-1)$ rounds, by Proposition 2.5. Therefore, since $\mathcal{V}$ is a vista taken at round $t^{\prime}$, all levels of $\mathcal{V}$ up to $L_{t^{\prime}-\tau(n-1)}$ are complete (recall that a level of a vista is complete if it has no missing nodes).
Assume Condition 2 first. Since $t^{\prime}\geq\tau(2n-c-1)+1$, all levels of $\mathcal{V}$ up to $L_{\tau(n-c)+1}$ are complete. Thus, level $L_{0}$ is complete, and therefore it has exactly $c$ nodes, because two agents are distinguishable at round 00 if and only if they have distinct inputs. Since the sum of the anonymities of these $c$ nodes is $n$, there must be fewer than $n-c$ branching nodes in $\mathcal{V}$ (excluding the root). Hence, the levels up to $L_{\tau(n-c)}$ contain an interval of at least $\tau$ consecutive levels, say from $L_{r}$ to $L_{r+\tau-1}$, where all nodes are non-branching and can be partitioned into $|L_{r}|=|L_{r+\tau-1}|$ strands $P_{1}$, $P_{2}$, …, $P_{|L_{r}|}$.
Note that a link between two agents at any round $r^{\prime}$ in the interval $[r+1,r+\tau]$ determines a pair of exposed nodes in $L_{r^{\prime}-1}$. Thus, by definition of $\tau$-union-connected network, the graph of exposed strands between $L_{r}$ and $L_{r+\tau-1}$ (constructed as $G$ in Line 12) is connected. It follows that the execution of Listing 1 terminates at Line 20 (as opposed to Line 6) whenever $t\geq r+\tau-1$. Thus, the procedure returns a pair $(t,S)$ with $0\leq t\leq r+\tau-1\leq\tau(n-c)$. In particular, all levels of $\mathcal{V}$ up to $L_{t+1}$ are complete.
@@
    $t^{\prime}\geq\tau(n-c)+d+1$, where $c$ is the number of distinct inputs held by agents at round 00,
@@ 4.2 Stabilizing Algorithm for Leaderless Networks
As a first application of Lemma 4.2, we will give stabilizing algorithm that efficiently computes the Input Frequency function $F_{IF}$ in all leaderless networks with a finite dynamic disconnectivity $\tau$, assuming no knowledge of $\tau$ or $n$. As a consequence, all frequency-based functions are efficiently computable as well, due to Proposition 2.2. Moreover, Proposition 6.3 states that no other functions are computable in leaderless networks, and Proposition 6.9 shows that our algorithm is asymptotically optimal.
@@
There is an algorithm that computes the Input Frequency function $F_{IF}$ on all networks in $\bigcup_{\tau\geq 1}\bigcup_{n\geq 1}\mathcal{N}^{\tau}_{n,0}$ and stabilizes in at most $\tau(2n-2)$ rounds.
@@
Our local algorithm is as follows. Each agent $p$ runs Listing 1 on its own vista $\mathcal{V}$, obtaining a pair $(t,S)$. If $t=-1$ or $S$ is not a homogeneous system of $k-1$ independent linear equations in $k$ variables, then $p$ outputs $\{(\lambda(p),1)\}$, where $\lambda(p)$ is its own input. Otherwise, since the rank of the coefficient matrix of $S$ is $k-1$, the general solution to $S$ has exactly one free parameter, due to the Rouché–Capelli theorem. Therefore, by Gaussian elimination, it is possible to express every variable $x_{i}$ as a rational multiple of $x_{1}$, i.e., $x_{i}=\alpha_{i}x_{1}$ for some $\alpha_{i}\in\mathbb{Q}^{+}$ (recall that the coefficients of $S$ are integers). Let $L_{t}=\{w_{1},w_{2},\dots,w_{k}\}$ and $L_{0}=\{v_{1},v_{2},\dots,v_{k^{\prime}}\}$. For every node $v_{i}\in L_{0}$, define $\beta_{i}\in\mathbb{Q}^{+}$ as $\beta_{i}=\sum_{w_{j}\in L_{t}\text{ descendant of }v_{i}}\alpha_{j}$, and let $\beta=\sum_{i}\beta_{i}$. Then, $p$ outputs
@@
The correctness and stabilization time of the above algorithm directly follow from Lemma 4.2. Specifically, if not all agents have the same input (hence $n\geq c\geq 2$), then at any round $t^{\prime}\geq\tau(2n-2)\geq\tau(2n-3)+1\geq\tau(2n-c-1)+1$, Condition 2 of Lemma 4.2 is met, and the system $S$ is satisfied by the anonymities of the nodes in $L_{t}$. Thus, $a(v_{i})=\alpha_{i}a(v_{1})$ for all $v_{i}\in L_{0}$, and therefore $\beta_{i}/\beta=a(v_{i})/n$. We conclude that, for any input assignment $\lambda$, the algorithm stabilizes on the correct output $\frac{1}{n}\cdot\mu_{\lambda}$ within $\tau(2n-2)$ rounds.
In the special case where all agents have the same input, then the default output $\{(\lambda(p),1)\}$ is correct since round 00. Also, whenever $S$ is a homogeneous system of $k-1$ independent linear equations, the algorithm still returns the correct output $\{(\lambda(p),1)\}$, because all nodes in level $L_{0}$ have the same label $\lambda(p)$. Hence, in this particular case the output is always correct. ∎
@@ 4.3 Terminating Algorithm for Leaderless Networks
For every $\tau\geq 1$ and $N\geq 1$, there is an algorithm that computes the Input Frequency function $F_{IF}$ on all networks in $\bigcup_{n\leq N}\mathcal{N}^{\tau}_{n,0}$ and terminates in at most $\tau(n+N-2)+1$ rounds.
@@
Given $\tau$ and $N$, our terminating local algorithm is as follows. Run Listing 1 on the agent’s vista $\mathcal{V}$, obtaining a pair $(t,S)$, and then do the same computations as in the algorithm for Theorem 4.4. If $t\geq 0$ and the current round $t^{\prime}$ satisfies $t^{\prime}\geq t+\tau(N-1)+1$, then the output is correct, and the agent terminates.
@@
We can also trade the knowledge of $\tau$ and $N$ for the knowledge of the dynamic diameter $d$ of the network (or an upper bound thereof).
@@
For every $d\geq 1$, there is an algorithm that computes the Input Frequency function $F_{IF}$ on all networks in $\bigcup_{n\geq 1}\mathcal{N}^{[d]}_{n,0}$ and terminates in at most $\tau(n-1)+d+1\leq dn+1$ rounds, where $\tau$ is the dynamic disconnectivity of the network.
@@
Note that knowledge of an upper bound $D\geq d$ is actually enough for the algorithm in Corollary 4.6 to work. Indeed, with the termination condition $t^{\prime}\geq t+D+1$, the same algorithm computes $F_{IF}$ on all networks in $\bigcup_{d\leq D}\bigcup_{n\geq 1}\mathcal{N}^{[d]}_{n,0}$ and terminates in at most $\tau(n-1)+D+1\leq d(n-1)+D+1$ rounds.